OpenStudy (anonymous):

Thomas needs to earn a B in his Geology class. His current test scores are 80, 82, 90, and 72. His final exam is worth 4 test scores. In order to earn a B, Thomas's average must lie between 80 and 89 inclusive. What range of scores can Thomas receive on the final exam to earn a B in the course? ***Please help!!! I need this fast and I also need the steps used to solve it. HELP!!!

OpenStudy (anonymous):

(80+82+90+72+4x)/5 = 80 (80+82+90+72+4y)/5 = 89 solve for x and y all possible answers are between them

OpenStudy (anonymous):

THANK YOU!!!! :D :D :D

OpenStudy (anonymous):

Oops I apologize don't divide by 5 divide by nine. I only considered the final as one grade instead of four. Sorry

OpenStudy (anonymous):

no its ok I still have to work it out.

OpenStudy (anonymous):

jpanct- Isn't it going to be a negative number???

OpenStudy (anonymous):

hmm... well it looks like if he wants an 80 average he will have to get a 99 on the final... are you sure that this question is valid?

OpenStudy (anonymous):

I'm not sure. I will talk to my teacher about it... Thanks I REALLY appreciate your hard work/help.

OpenStudy (anonymous):

(80+82+90+72+4x)/8 = 80 (80+82+90+72+4y)/8 = 89 x= 79 y= 97 he needs between a 79 and 97 for an average between 80 and 89. Sorry for the previous screwups idk what i was doing

OpenStudy (anonymous):

Don't worry about it. we both kinda messed it up.